# Event2Vec: Processing Neuromorphic Events directly by Representations in Vector Space

### v3

@@ A.3 Model Hyper-parameters
Unless otherwise stated, all models were trained using BFloat16 mixed precision. The training configuration for all models includes a base learning rate of $lr_{b}=0.001$, a batch size of 64, and the AdamW optimizer (Loshchilov & Hutter, 2019), conducted over 64 epochs. The effective learning rate is determined by a linear scaling rule based on the number of GPUs ($n_{gpus}$) used in distributed data-parallel training: $lr=lr_{b}\cdot n_{gpus}/256$. A warmup phase is implemented for the first four epochs, during which the learning rate is linearly increased from $0.01\cdot lr$ to its full value, $lr$. For the subsequent epochs, a cosine annealing schedule (Loshchilov & Hutter, 2017) is employed to gradually reduce the learning rate to a minimum value, $lr_{min}$. For the DVS Gesture and ASL-DVS datasets, both weight decay and label smoothing were disabled. In contrast, for the DVS-Lip classification task, we set the weight decay to 0.05 and applied label smoothing with a factor of 0.1.
@@
Table 5 provides a detailed summary of the model-specific hyper-parameters. Here, $D$ denotes the embedding dimension, $D_{f}$ represents the hidden feature dimension of the feed-forward neural network (FFN), and $n_{head}$ is the total number of attention heads. The repeats parameter specifies how many times the training set is iterated through within a single epoch. Notably, the number of heads for the key (k) and value (v) projections is set to $n_{head}/2$, and group normalization (Wu & He, 2018) is applied to both. To prevent exploding gradients, we employ gradient clipping, capping the $L_{2}$ norm of the gradients at 1.0.
@@
Table 5: Hyper-parameters of training models for classification tasks on different datasets
Dataset $D$ $D_{f}$ $n_{head}$ Depth Repeats $n_{gpus}$ $lr_{min}$
@@
A.4 Data augmentations
@@
Denote $\mathcal{U}(a,b)$ as the uniform distribution between $a$ and $b$, and $\text{RandInt}(m,n)$ as a random integer taken from the set m,m+1,,n{m,m+1,...,n}, where each integer has an equal probability of being selected.
@@
For an event stream, the data augmentations are applied on events directly. For simplicity, we omit the event index in this subsection. Unless otherwise specified, augmentations are applied on each event stream independently. Note that the coordinates are converted to floating percision before applying any augmentation. After all augmentations are applied, coordinates will be quantized, and only events whose coordinates are valid, i.e., $x\in[0,W-1],y\in[0,H-1]$, are kept.
@@
  • Random Erasing: Erase an $h\times w$ area with $h,w\sim\mathcal{U}(0,16)$ with the probability 0.1. The center of this area $(c_{x},c_{y})$ satisfy $c_{x}\sim\mathcal{U}(0,W-1),c_{y}\sim\mathcal{U}(0,H-1)$.
@@
  • Random Resizing: Resize $(x,y)$ to $(s_{x}\cdot x,s_{y}\cdot y)$ where $s_{x},s_{y}\sim\mathcal{U}(0.8,1.2)$.
@@
  • Random Flip: The event stream is flipped horizontally with a probability of $0.5$.
@@
The augmentations listed above are each applied independently with a probability of $0.5$. The token-mix is applied on the embedding tensor with probability $0.5$. Specifically, when training on cluster events, the intensity $\rho$ is randomly set to $1$ with a probability of $0.1$. We use drop path (Larsson et al., 2016) in the linear attention layer, with the probability increasing linearly from 0 to $0.4$ with depth.
A.5 Self-supervised Training Details
A binary mask m of shape $(B,L)$ is generated from a Bernoulli distribution. The probability of masking any given token is set to 0.3, which defines the mask ratio. Each token $\textbf{v}[i][j]$ corresponding to a mask entry $\textbf{m}[i][j]=1$ is replaced by a single, learnable, $D$-dimensional mask token $\textbf{v}_{m}$. This operation results in a corrupted embedding tensor, denoted as $\hat{\textbf{v}}$. Concurrently, the original coordinates $(\textbf{x}_{m},\textbf{y}_{m},\textbf{p}_{m})$ of the masked tokens are preserved to serve as the ground truth for the reconstruction loss.
The corrupted tensor $\hat{\textbf{v}}$ is then processed by the model’s linear attention layers. Following this, the output embeddings that correspond to the initially masked positions, denoted $\hat{\textbf{v}}_{m}$, are extracted from the final output tensor using the mask m.
$\hat{\textbf{v}}_{s}=\frac{\hat{\textbf{v}}_{m}}{\bm{\rho}}-\textbf{v}_{t}.$ (14)
[…]
A.6 Experimental Details for Pre-processing Latency
The pre-processing latency benchmarks, with results presented in Table 2, were conducted on a Red Hat Enterprise Linux 8.10 server. This server was equipped with an NVIDIA A100 GPU (80GB PCIe), an Intel Xeon Gold 6326 CPU (utilizing 8 cores), and 256GB of RAM. To mitigate the impact of data I/O, the DVS Gesture dataset was loaded entirely into RAM for the duration of the experiments. For each method, we optimized the batch size and the number of workers to achieve the minimum possible latency. All implementations were based on the publicly available source code from their respective original publications or other high-performance software libraries. The reported latency for each method is the average of 8 measurement runs, which were preceded by 2 warm-up runs to ensure system stability. Further implementation details are provided in Table 6.
Table 6: Implementation details of Table 2
Method Batch size Workers Software Library
Random Sampling 64 64 Numpy (Harris et al., 2020b)
Frame 8 8 SpikingJelly (Fang et al., 2023)
Graph 8 8 PyTorch Geometric (Fey & Lenssen, 2019)
Voxel Grid 16 8 Tonic (Lenz et al., 2021)
K-Means 2 2 Faiss (Johnson et al., 2019)
Window Slicing 8 8 Original paper (Santambrogio et al., 2024)
A.7 Experimental Details for Model Throughput and Latency
The throughput and latency experiments, with results presented in Table 3, were conducted under the same operating system and hardware environment as the pre-processing benchmarks, which are detailed in Appendix A.6. The dataset was also pre-loaded into RAM to eliminate I/O bottlenecks.
For each model, we performed a search for the optimal batch size to maximize performance. The optimal batch sizes were determined to be 64 for Sparse GRU, 16 for FARSE-CNN, 64 for AEGNN, 64 for SNN and 512 for event2vec, respectively. We observed that a fixed number of 8 workers yielded the best performance across all models. The benchmarking process for each model involved an initial warm-up phase, followed by multiple measurement runs, the results of which were then averaged. Due to significant variations in computational cost among the models, the number of warm-up iterations and measurement batches was tailored for each specific model. However, we ensured that the number of batches was sufficiently large, such that further increases did not yield any significant changes in the measured performance, confirming the stability of our results.
Regarding the implementations, for Sparse GRU [^1] https://github.com/Efficient-Scalable-Machine-Learning/EvNN and FARSE-CNN [^2] https://github.com/AIRLab-POLIMI/farse-cnn, the complete source code was available in the official GitHub repositories, and we used them directly. For AEGNN, only the inference code was publicly available [^3] https://github.com/uzh-rpg/aegnn. We therefore implemented the necessary Dataset class and model architecture by referencing their provided implementation for the N-Cars dataset (Sironi et al., 2018). For SNNs, the model used by Yao et al. (2023) enhances the Parametric Leaky Integrate-and-Fire Neuron Network (PLIF-Net) (Fang et al., 2021) with attentions, but the source code is not released. Given these additional attention modules only add slight complexity, we evaluate on the PLIF-Net as an alternative. SpikingJelly (Fang et al., 2023) provides a high-performance implementation [^4] https://spikingjelly.readthedocs.io/zh-cn/latest/activation_based_en/classify_dvsg.html for the PLIF-Net with advanced accelerating techniques, and we benchmark on code from SpikingJelly directly.
All benchmarks were running in BFloat16 mixed precision except for Sparse GRU, which depends on the Haste library (Nanavati, 2020) with only supports Float32.
A.8 Experimental Details for the Ablation Study on Event Numbers
The ablation study on the number of events, with results reported in Figure 3, evaluates the impact on model throughput, latency, and accuracy. These experiments were conducted using the same operating system and hardware environment detailed in Appendix A.6. During the training process on the DVS Gesture dataset, the chunk length parameter $l_{chunk}$ for the temporal chunk dropout augmentation is scaled proportionally with the number of events. Furthermore, for experiments using 64 and 32 events, all data augmentation techniques are disabled. This measure is implemented to prevent the augmentation from inadvertently removing all events, which would subsequently lead to a NaN (Not-a-Number) loss.
A.9 Visualization of Neighborhood Semantics
Figure 8: Visualization of the parametric embedding weight $\textbf{W}_{\phi}$ and the standard embedding weight $\textbf{W}_{s}$ in the vector field.

### v4

@@ A.2 Model Structures and Hyper-parameters
Unless otherwise stated, all models were trained using BFloat16 mixed precision. The training configuration for all models includes a base learning rate of $lr_{b}=0.001$, a batch size of 64, and the AdamW optimizer (Loshchilov and Hutter, 2019) for 64 epochs. The effective learning rate is determined by a linear scaling rule based on the number of GPUs ($n_{gpus}$) used in distributed data-parallel training: $lr=lr_{b}\cdot n_{gpus}/256$. A warmup phase is implemented for the first 4 epochs, during which the learning rate is linearly increased from $0.01\cdot lr$ to $lr$. For the subsequent epochs, a cosine annealing schedule (Loshchilov and Hutter, 2017) is employed to gradually reduce the learning rate to a minimum value, $lr_{min}$. For the DVS Gesture and ASL-DVS datasets, both weight decay and label smoothing were disabled. In contrast, for the DVS-Lip classification task, we set the weight decay to 0.05 and applied label smoothing with a factor of 0.1.
@@
Figure 6 illustrates the detailed network architecture, including (a) the spatial embedding architecture of event2vec, (b) the temporal embedding architecture of event2vec, and (c) the architecture of the entire network. Table 5 provides a detailed summary of the model-specific hyper-parameters. Here, $D$ denotes the embedding dimension, $l$ is the number of Transformer blocks in Backbone, $D_{f}$ represents the hidden feature dimension of the feed-forward neural network (FFN), and $n_{head}$ is the total number of attention heads. The repeats parameter specifies how many times the training set is iterated through within a single epoch. Notably, the number of heads for the key (k) and value (v) projections is set to $n_{head}/2$, and group normalization (Wu and He, 2018) is applied to both. To prevent exploding gradients, we employ gradient clipping, capping the $L_{2}$ norm of the gradients at 1.0.
@@
Table 5: Hyper-parameters of training models for classification tasks on different datasets.
Dataset $D$ $D_{f}$ $n_{head}$ $l$ Repeats $n_{gpus}$ $lr_{min}$
@@
A.3 Accuracy with Different Types of Self-Attention
In addition to the Forgetting Transformer (FoX), we also evaluated the performance when using Gated Linear Attention (GLA), with the results presented in Table 6. The results show that GLA achieves high performance on relatively easy classification tasks such as DVS Gesture and ASL-DVS; however, its performance drops by 3.53% on the more challenging DVS-Lip classification task, which may be attributed to the fact that linear attention struggles to prevent the decay of long-term memory when processing long input sequences. Overall, the performance degradation across all three tasks is negligible when replacing FoX with GLA, indicating that Event2Vec is not sensitive to the specific type of attention mechanism adopted.
Table 6: Comparison of accuracy with different types of self-attentions.
Dataset Accuracy of FoX (%) Accuracy of GLA (%)
[…]
DVS-Lip 75.88 72.35
A.4 Bi-directional Self-Attentions
We evaluated two variants of Self-Attention in our experiments, namely Gated Linear Attention (GLA) (Yang et al., 2024) and Forgetting Transformer (FoX) (Lin et al., 2025). GLA is a typical linear attention mechanism, which can be regarded as a special case of Recurrent Neural Networks (RNNs) (Katharopoulos et al., 2020), where the input sequence order affects the output results. Although FoX does not fall into the category of linear attention, it adopts an RNN-style gating mechanism that depends on input sequence order, thus the input order also exerts an impact on its outputs. Due to the fixed and limited size of hidden states, RNNs inevitably suffer from long-distance information attenuation when processing ultra-long sequences. To mitigate the degradation of long-term memory, we extend GLA and FoX to bidirectional variants.
We adapt this formulation to be bi-directional by inputting both forward and reversed $\textbf{Q},\textbf{K},\textbf{V}$. The bi-directional outputs are computed as:
$\textbf{O}_{f}$ $=\text{Attention}(\textbf{Q},\textbf{K},\textbf{V}),$ (7)
$\textbf{O}_{b}$ $=\text{Reverse}(\text{Attention}(\overleftarrow{\textbf{Q}},\overleftarrow{\textbf{K}},\overleftarrow{\textbf{V}})),$ (8)
$\textbf{O}_{fb}[t]$ $=\textbf{W}_{fb}[\textbf{O}_{f}[t];\textbf{O}_{b}[t]].$ (9)
Unlike classic bi-directional RNNs (Schuster and Paliwal, 1997) that often use independent parameters for each direction, our model employs shared parameters for the two directions. By sharing the projection weights ($\textbf{W}_{q},\textbf{W}_{k},\textbf{W}_{v},\textbf{W}_{forget}$) across both passes, we ensure that the parameter count remains comparable to the uni-directional baseline, with the only increase arising from the fused output projection $\textbf{W}_{fb}$.
We also tested the performance changes when using bi-directional self-attentions without parameter sharing, and the results are summarized in Table 7. The results show that the number of parameters increases by approximately 25% when parameters are not shared. Although the fitting capacity is theoretically improved without parameter sharing, the test-set accuracy instead decreases, indicating slight overfitting. This experimental result demonstrates that our bi-directional self-attentions with parameter sharing not only reduce the number of parameters but also mitigate overfitting.
Table 7: Changes in accuracy and parameters when using bi-directional self-attentions without parameter sharing.
Dataset Parameters (MB) Accuracy (%)
DVS Gesture 0.65 (+25%) 96.63 (-0.94)
ASL-DVS 0.34 (+26%) 99.86 (-0.05)
DVS-Lip 22.90 (+25%) 75.36 (-0.52)
A.5 Self-supervised Training Details
A binary mask m of shape $(B,L)$ is generated from a Bernoulli distribution. The probability of masking any given token is set to $30\%$, which defines the mask ratio. To prevent the model from making predictions via simple interpolation, we mask out $l_{mask}$ consecutive tokens in sequence. The length of each masked span $l_{mask}$ is sampled from a geometric distribution $l_{mask}\sim\text{Geometric}(p)$ with $p=0.1$, resulting in an average length of $10$. Each token $\textbf{v}[i][j]$ corresponding to a mask entry $\textbf{m}[i][j]=1$ is replaced by a single, learnable, $D$-dimensional mask token $\textbf{v}_{m}$. This operation results in a corrupted embedding tensor, denoted as $\hat{\textbf{v}}$. Concurrently, the original coordinates $(\textbf{x}_{m},\textbf{y}_{m},\textbf{p}_{m})$ of the masked tokens are preserved to serve as the ground truth for the reconstruction loss.
The corrupted tensor $\hat{\textbf{v}}$ is then processed by the model’s linear attention layers. Following this, the output embeddings that correspond to the initially masked positions, denoted as $\hat{\textbf{v}}_{m}$ are extracted from the final output tensor using the mask m.
$\hat{\textbf{v}}_{s}=\frac{\hat{\textbf{v}}_{m}}{\bm{\rho}}-\textbf{v}_{t}.$ (10)
[…]
A.6 Experimental Environment for Performance Testing
All experiments related to performance testing (such as throughput and latency) mentioned in this paper were conducted on a Red Hat Enterprise Linux 8.10 server. This server was equipped with an NVIDIA A100 GPU (80GB PCIe), an Intel Xeon Gold 6326 CPU (utilizing 8 cores), and 256GB of RAM. To mitigate the impact of data I/O, all datasets were loaded entirely into RAM for the duration of the experiments.
The three models included for comparison in Table 2 are Max-Former (Fang et al., 2025), GNN & Transformer (Yuan et al., 2023), and Spiking ResNet18 & BiGRU (Dampfhoffer and Mesquida, 2024). All of these models provide official open-source code, which allowed us to conduct experiments based on their codebases.
A.7 Impact of the number of events on DVS Gesture
As illustrated in Figure 7, we benchmark the impact of varying the number of randomly sampled events ($L$) on several key metrics: training/inference throughput, single event stream inference latency, and accuracy on the DVS Gesture dataset. Experimental results on the ASL-DVS dataset show similar trends, and thus are not presented here. As $L$ increases, both training and inference throughput drop sharply, while the single-stream latency remains nearly unchanged. This indicates that for a single sample, the dominant latency arises from the CUDA kernel launch overhead rather than computation itself. Meanwhile, the accuracy improves rapidly with the growth of $L$, demonstrating that more events facilitate the model’s decision-making process.
Figure 7: Effect of number of events on DVS Gesture.
A.8 Visualization of Neighborhood Semantics
Figure 9: Visualization of the parametric embedding weights $\textbf{W}_{\phi}$ and the standard embedding weight $\textbf{W}_{s}$ as a vector field.
A.9 Data Augmentations
Transformers (including linear attention) lack inductive bias and thus require more data for learning. We used data augmentation methods to expand the data volume to a certain extent, thereby improving performance. Specifically, we did not use data augmentation on the ASL-DVS dataset because we found that SOTA (State-of-the-Art) performance could be achieved without it—this is likely due to the sufficient scale of this dataset: the number of samples in its training set is approximately 80,640, while that of DVS Gesture is 1,176, and DVS-Lip is 14,896.
@@ A.9 Data Augmentations
Denote $\mathcal{U}(a,b)$ as the uniform distribution between $a$ and $b$, and $\text{RandInt}(m,n)$ as a random integer taken from the set $\{m,m+1,...,n\}$, where each integer has an equal probability of being selected.
@@
For an event stream, the data augmentations are applied on events directly. For simplicity, we omit the event index in this subsection. Unless otherwise specified, augmentations are applied on each event stream independently. Note that the coordinates are converted to floating-point percision before applying any augmentation. After all augmentations are applied, only events whose coordinates are valid, i.e., $x\in[0,W-1],y\in[0,H-1]$, are kept.
@@
  • Random Erasing: Erase an $h\times w$ area with $h,w\sim\mathcal{U}(0,16)$ with a probability of 0.1. The center of this area $(c_{x},c_{y})$ satisfy $c_{x}\sim\mathcal{U}(0,W-1),c_{y}\sim\mathcal{U}(0,H-1)$.
@@
  • Random Flip: The event stream is flipped horizontally with a probability of $1$.
@@
  • Random Chunk Drop: Randomly mask $n$ temporal chunks with $n=4$. The length of each chunk $l$ is sampled from $\mathcal{U}(1,128)$ and scaled by the ratio of valid events. The starting position $s$ satisfies $s\sim\mathcal{U}(0,L-1)$.
@@
The augmentations listed above are each applied independently with a probability of $0.5$. The token-mix is applied on the embedding tensor with a probability of $0.5$. Specifically, when training on cluster events, the intensity $\rho$ is randomly set to $1$ with a probability of $0.1$. We use drop path (Larsson et al., 2016) in the linear attention layer, with the probability increasing linearly from 0 to $0.4$ with depth.
In addition, all other methods in Tab 1 also use data augmentation, which is summarized in Tab 8.
Table 8: Data augmentation strategies for methods in Table 1.
Dataset Method + Representation Data Augmentations
DVS Gesture Sparse GRU + Frame (Subramoney et al., 2023) Random crop, translation, and rotation
SNN + Frame (Yao et al., 2023) Random slice and integrate
FARSE-CNN + Window Slicing (Santambrogio et al., 2024) Random coordinate translations
Event MAE + Point Cloud (Sun et al., 2025) Point resampling from Point-BERT (Yu et al., 2022)
Max-Former + Frame (Fang et al., 2025) Mixup and Cutmix
Linear Attention + Event2Vec Random resize, rotation, shear, translate, erase, and chunk dropout
ASL-DVS GNN,CNN + Graph (Bi et al., 2019) Random scale, flip, and rotation of node positions
GNN & Transformer + Image & Voxel Graph (Yuan et al., 2023) Random scale and translate
Linear Attention + Event2Vec None
DVS-Lip ResNet-18,BiGRU + Frame (Tan et al., 2022) Random crop and horizontal flip
Spiking ResNet18 & BiGRU + Frame (Dampfhoffer and Mesquida, 2024) Random crop, horizontal flip, spatial masking, zoom, and temporal mask
Linear Attention + Event2Vec Random resize, rotate, shear, flip, translate, and erase